If the equations x^2+ ax + b =0 and x^2 + bx +a =0 have exactly one common root, then numerical value of a+ b is

Grade:

1 Answers

AJIT AskiitiansExpert-IITD
13 years ago

Dear sudha ,

let common root be k

K^2 + ka +b =0

K^2 + kb +a =0

subtract the equations

k = 1     putting this value

1 +a+b =0   i.e  a+b =-1

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Win exciting gifts by answering the questions on Discussion Forum..